Solve by addition.
-3x + y = 3
2x - 3y = 5

Answer by checkley71(8403)

MULTIPLY EQUATION #1 BY 3 WE GET -9X+3Y=9 NOW ADD THIS TO EQUATION #2 WE GET
2X-3Y=5
-9X+3Y=9
---------------
-7X=14 OR X=14/-7 OR X=-2
THEN 2(-2)-3Y=5 OR -4-3Y=5 OR -3Y=9 OR Y=9/-3 OR Y=-3